OpenStudy (anonymous):

I have to find the vertex of: y= -3x^2+12x-8 and I have to plot it on a graph, providing an ordered pair as the vertex. I'm not understanding this at all. How should I begin?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

complete the sqaure

OpenStudy (anonymous):

first factor out x^2 coeffiecnet

OpenStudy (anonymous):

y = -3 ( x^2 -4x ) -8

OpenStudy (anonymous):

note, we only factored out common factors on the x terms

OpenStudy (anonymous):

then complete the sqaure in side the brackets

OpenStudy (anonymous):

ok, it kinda makes sense so far..

OpenStudy (anonymous):

to complete the square on x^2 -4x , we need to half the -4 ( to get -2 ) , the square it , and add and subtract inside the brackets

OpenStudy (anonymous):

so y= -3 ( x^2 -4x + 4 - 4 ) - 8 // notice, its important you add and substract INSIDE THE BRACKET

OpenStudy (anonymous):

we need to maintain equality

OpenStudy (anonymous):

alright, with you on that

OpenStudy (anonymous):

now you group the perfect sqaure inside the bracket

OpenStudy (anonymous):

so y= -3 ( (x-2)^2 -4) -8

OpenStudy (anonymous):

now you expand the brackets , being careful with the signs

OpenStudy (anonymous):

so y= -3(x-2)^2 +12 -8

OpenStudy (anonymous):

So I'm seeing here the only way the vertex can be solved is by doing the equation, then plotting it on the graph- how would I plot this #1, and #2, how would I come up with a solution set??

OpenStudy (anonymous):

y= -3(x-2)^2 +4

OpenStudy (anonymous):

vertex = (2,4)

OpenStudy (anonymous):

you could do it other ways , you might the equation of axis of symmetry of a parabola is x=-b/2a

OpenStudy (anonymous):

And how to plot? I have to do it on an online graph. Totally confused with it

OpenStudy (anonymous):

and you know the vertex is on the axis of symmetry

OpenStudy (anonymous):

I dont know there, I know how to do a general plot but have idea how the computer with mark you

OpenStudy (anonymous):

"graph the equation by hand; find the vertex and round approximate coordinates to the second decimal place."

OpenStudy (anonymous):

"Use the graphing tool to graph the parabola"

OpenStudy (anonymous):

THIS is what is confusing me. Do I graph the solution set, I'm just not getting it. I understand for the most part the equation, but not plotting

OpenStudy (anonymous):

One more equation- (x+8)^2+8(x+7)=64. I can solve by the method of my choice

OpenStudy (anonymous):

Its not a negative sign in front of the x, I just need a solution set